# DeepEdit: Knowledge Editing as Decoding with Constraints

### v3

@@ Abstract
Answering multi-hop questions involving new knowledge is a challenging task in evaluating large language models’ (LLMs) knowledge editing (KE) methods. This task is rather difficult because the LLMs’ hallucinations on new knowledge would harm the logical coherence of LLMs’ multi-hop reasoning and lead to incorrect answers. To address this issue, we design decoding constraints to “regulate” LLMs’ reasoning, enhancing logical coherence when incorporating new knowledge. We incorporate the constraints into a new KE framework: DeepEdit ( Depth-first Search-based Constrained Decoding for Knowledge Editing), which enhances LLMs to generate coherent reasoning chains with new knowledge through a depth-first search. Our search selects the most important knowledge that satisfies our constraints as the reasoning step to efficiently increase the reasoning depth. In addition to DeepEdit, we propose two new KE benchmarks: MQuAKE-2002 and MQuAKE-hard, which provide more precise and challenging assessments of KE approaches. Qualitatively, DeepEdit enables LLMs to produce succinct and coherent reasoning chains involving new knowledge. Quantitatively, it yields significant improvements on multiple KE benchmarks.
@@ 1 Introduction
Knowledge editing (KE) aims to enable large language models (LLMs) to produce outputs in accordance with the knowledge that is new or conflicted with the knowledge memorized by the LLMs. A challenging task to evaluate KE is the multi-hop question answering with new knowledge Zhong et al. (2023). In this task, LLMs should answer the questions corresponding to a chain of facts including new knowledge (see Figure 1(a)).
@@
Recent literature has mainly sought to perform KE through in-context editing Zhong et al. (2023); Cohen et al. (2023); Zheng et al. (2023), which puts new knowledge in the input context to fulfill KE. However, the input-conflicting hallucinations on the in-context new knowledge would harm the logical coherence of multi-step reasoning and lead to incorrect answersZhang et al. (2023).
Taking a closer look at the ground-truth output in Figure 1(a), we notice that: the adjacent reasoning steps should be logically coherent, and new knowledge should participate in the reasoning chain at appropriate positions. This raises a natural question:
Can we directly control the decoding of LLMs’ to soundly incorporate the new knowledge?
@@
The central idea of this paper is to view KE as a problem of constrained decoding. We design the decoding constraints, by following which LLMs can soundly incorporate new knowledge in LLMs reasoning. Specifically, we propose the following constraints: Conciseness requires every reasoning step to be identical; Coherence requires the adjacent reasoning steps to be coherent; Receptiveness requires new knowledge to replace the conflicted memorized knowledge; Pertinence requires every reasoning step to be relevant to the target question (see Figure 2). A multi-step reasoning chain that leads to the correct answer should satisfy the above constraints.
@@
We incorporate the above constraints into a new KE framework: DeepEdit ( Depth-first Search-based Constrained Decoding for Knowledge Editing). When producing every reasoning step, DeepEdit searches over the parametric and new knowledge to find the ones that satisfy our constraints, termed as valid step candidates. Among the valid candidates, DeepEdit takes the most important one as the new reasoning step for further reasoning, which is the (new) knowledge that is most helpful to answer the target question. Our depth-first search strategy efficiently increases the reasoning depth and enhances LLMs to produce the coherent reasoning chains involving new knowledge.
The advantages of our DeepEdit on KE are three-folds. First, it does not need to retrain the LLMs like prior model editing methods Zhu et al. (2020), which is complex and expensive. Our DeepEdit works during inference with negligible costs. Second, our DeepEdit is flexibly applicable to any black-box LLM without requiring access to model parameters or token-wise distributions. Third, our DeepEdit enhances LLMs to soundly utilize new knowledge to produce coherent reasoning chains in accordance with new knowledge.
@@
To provide more precise and challenging assessments of KE methods, we build two new benchmarks MQuAKE-2002, and MQuAKE-hard, which resolve the knowledge-conflicting annotation mistakes in the popular KE benchmark MQuAKE-3k Zhong et al. (2023). We evaluate DeepEdit on the KE benchmarks MQuAKE-3k, MQuAKE-2002, and MQuAKE-hard Zhong et al. (2023); Meng et al. (2022a). Qualitatively, DeepEdit enhances LLMs to produce more succinct and faithful reasoning with new knowledge, as shown in Figure 9. Quantitatively, DeepEdit improves the KE for popular LLMs by a significant margin OpenAI (2022); Touvron et al. (2023).
@@ 2 Related Work
Previous studies have explored multiple methods for editing the knowledge of LLMs by introducing new knowledge into static model artifacts Zhu et al. (2020); Sotoudeh and Thakur (2019); Dai et al. (2021); Hase et al. (2021); Zhou et al. (2023); Dong et al. (2022); Huang et al. (2023). Some of these approaches involve identifying and modifying model weights associated with specific concepts Meng et al. (2022a, b); Dai et al. (2021), as well as rapid adaptation facilitated by a compact auxiliary editing network Mitchell et al. (2022a); De Cao et al. (2021); Mitchell et al. (2021). Given the fast-growing parameter sizes of large language models (LLMs) Zhao et al. (2023), frequently updating LLMs with new knowledge through retraining is more and more expensive. Hence, it is vital to effectively edit the LLMs’ knowledge without retraining.
Recent work observes the superior KE performance of in-context editing. MeLLo Zhong et al. (2023) designs a single prompt to alternately conduct the text generation and new knowledge analysis using the LLM generator. Cohen et al. (2023) propose to append the new knowledge the head of the input prompt. In-context editing lacks direct control of the LLMs’ outputs and how the new knowledge influences the outputs is not traceable. Different from them, we propose the new decoding method DeepEdit for KE, which effectively augments LLMs on KE during the decoding stage. We illustrate a comparison between DeepEdit and existing KE methods in Table 1.
Specifically, in our work, we consider a challenging setting of knowledge editing: the multi-hop question answering (QA) with knowledge editing Zhong et al. (2023). In-context learning methods combined with an optional retrieval module have become a popular approach in handling multi-hop QA tasks Press et al. (2022); Yao et al. (2022); Khattab et al. (2022); Mitchell et al. (2022b). We propose four types of decoding constraints to enhance question answering with KE. Our work improves multi-hop QA with KE by applying constraints to the reasoning of LLMs, which directly controls LLMs’ outputs to follow new knowledge.
@@ 3 Methodology
Multi-hop question answering with new knowledge is a challenging task for assessing knowledge editing (KE) of large language models (LLMs) Zhong et al. (2023). This task considers the questions that correspond to a chain of facts including new knowledge. It assumes a database that stores all new facts, which LLMs should adhere to, and posits that only a few facts can influence the ground-truth answer to each question. In this section, we will introduce how we integrate the decoding constraints into the design of a novel KE method, DeepEdit (Depth-first Search based Decoding for Knowledge Editing), to enhance the KE of black-box LLMs.
@@ 3.1 Constraint Designing for Knowledge Editing
On a LLM’s multi-step reasoning chain to answer a multi-hop question involving new knowledge, we find the following constraints on reasoning steps can help LLMs to give the correct answer:
@@
Among the above four constraints, Conciseness removes loops in the reasoning chain; Coherence guarantees the coherence of reasoning; Receptiveness ensures the LLMs’ receptiveness to new knowledge; Pertinence guides LLMs to answer the target question. We visualize different constraints in Figure 2. Intuitively, the above constraints are easy to follow for LLMs when there is no new knowledge. However, the LLMs’ utilization of new knowledge is affected by the hallucination, which makes the above constraints easy to break when incorporating new knowledge. In this work, we apply the above constraints to incorporate the new knowledge into LLMs’ decoding.
@@ 3.2 Constraint Verification on Every Step
We term the reasoning steps that satisfy all the constraints as valid steps. Because an invalid reasoning step could mislead LLMs into generating incorrect subsequent steps, we verify the constraints during the decoding of each reasoning step.
At every iteration, we take both the parametric and new knowledge as the step candidates. First, we let LLMs generate one reasoning step with temperature as 00, which represents the parametric knowledge that the LLM believes to be useful to answer the target question. Second, we retrieve $N$ new knowledge closest to the generated reasoning step. The distance between knowledge is measured by the semantic distance of sentence-level embeddings given by a pre-trained BERT model Devlin et al. (2019). The closer the context of the new knowledge is to the generated step, the more likely it is that the new knowledge will be helpful in answering the target question. We take these $N+1$ knowledge as the reasoning step candidates, with examples shown in Figure 3.
We verify the constraints on the above step candidates to find the valid step candidates. In principle, every constraint’s verification can be seen as a binary classification problem on a pair of sentences. For example, Coherence’s verification can be completed by assessing whether a pair of reasoning steps is relevant to each other or not. Therefore, one option to do constraint verification is to train a binary classifier for every constraint. However, such a solution is not data or time-efficient. In this work, we develop four specialized in-context learning-based LLM agents to act as the verifiers. Every verifier’s prompt includes $D$ positive and negative demonstrations randomly sampled from the instances of the MQuAKE dataset Zhong et al. (2023). In our experiments, we set $D=2$ and $N=2$ by default.
@@ 3.3 Efficient Reasoning with Depth-first Search
At every iteration, there may exist more than valid step candidates. We can take every one of them as the new reasoning step in parallel and make further reasoning based on each of them separately. This design can be seen as a breadth-first search of the LLMs’ multi-hop reasoning chain incorporating new knowledge. It is undoubtedly time-consuming because the time complexity grows exponentially with the number of reasoning steps. This raises a natural question:
Can we more efficiently expand the depth of reasoning chains during decoding?
Our answer is yes. We propose to take the most important step candidate as the reasoning step for further reasoning. Actually, different valid step candidates are not equally important. First, the edited facts should be more important than the generated step, since the former generally contains new knowledge unknown by LLMs. Second, the edited facts that are closer to the generated step are more important, which are more helpful in answering the target question due to their relevance to the LLM’s generated reasoning context. Based on the above prior, in every iteration, we select the most important knowledge as the next reasoning step and do further reasoning based on it, as shown in Figure 4.
@@
We name our method as DeepEdit: depth-first search-based constrained decoding for knowledge editing. DeepEdit takes the most important step candidate as the new reasoning step at every iteration and efficiently produces the multi-hop reasoning incorporating new knowledge. We summarize DeepEdit in Algorithm 1 and visualize it in Figure 1. Furthermore, we present the iteration-wise outputs of our DeepEdit in Figure 8.
Assume the number of reasoning hops as $h$, and the number of valid step candidates per iteration as $c$. The time complexity of breadth-first search is $c^{h}$. In contrast, our DeepEdit has a much lower time complexity of $\mathcal{O}(h)$ thanks to our depth-first search design considering the importance of different candidates. Empirical results validate DeepEdit’s superior efficiency compared with the breadth-first search.
The strategy of depth-first search has been utilized by some research on LLMs, such as tree-of-thoughts Yao et al. (2023). Different from them, our DeepEdit’s depth-first search is designed to efficiently find the multi-hop reasoning chain incorporating new knowledge. Our DeepEdit is free to be applied to all black-box LLMs: it bypasses the requirement of accessing the token-wise distributions because all the decoding actions are defined at the step level.
Algorithm 1 DeepEdit: Depth-first Search based Constrained Decoding for Knowledge Editing.
Input: The query question $q$. The prompt $p$. The set of new knowledge $\mathcal{K}$. The number of retrieved new knowledge $N$. Maximum number of reasoning depth $r$.
Output: The answer $a$ to the query question.
1:for $i\leftarrow 1$ to $r$ do
2:     $o\leftarrow\mbox{LLM}(p)$
3:     if $o$ is the answer then
4:         return $o$
5:     $\hat{\mathcal{K}}\leftarrow\mbox{search}(o,\mathcal{K},n)$triangleright retrieve the $n$ semantically similar new knowledge to $o$, ranked by the distance to $o$ in the ascending order.
6:     for $j\leftarrow 1$ to $n$ do
7:         $k\leftarrow\hat{\mathcal{K}}[i]$
8:         $nk\leftarrow\mbox{False}$triangleright $nk$ is the boolean variable to decide whether some new knowledge is valid.
9:         if ($k$ satisfies all the constraints) then
10:              $nk\leftarrow\mbox{True}$
11:              $p\leftarrow p+k$
12:              break
13:     if not $nk$ then
14:         $p\leftarrow p+o$      return $o$
@@ 5.1 Experimental Settings
LLM as Constraint Verfier Conciseness Coherence Receptiveness Pertinence LLaMA2-7B-Chat 82.3 77.9 86.3 78.2 GPT-3.5-Turbo-Instruct 92.8 91.5 92.2 86.7
@@ Appendix A Analysis on Knowledge Editing Batch Size
Figure 8: The iteration-wise visualization of our DeepEdit on an instance in MQuAKE-3k. The chosen step candidate in every iteration that satisfies all the constraints are in green. The dropped step candidates are followed by the constraints that they do not satisfy.
@@ Appendix B Analysis on the Data where the Edited Facts do not Change the Ground-Truth Answers
Appendix C Iteration-wise Visualization of Our DeepEdit
We present the iteration-wise outputs of our DeepEdit on an instance in Figure 8.
Appendix D Case Study

### v4

@@ Abstract
How to edit the knowledge in multi-step reasoning has become the major challenge in the knowledge editing (KE) of large language models (LLMs). The difficulty arises because the hallucinations of LLMs during multi-step reasoning often lead to incorrect use of new knowledge and incorrect answers. To address this issue, we design decoding constraints to “regulate” LLMs’ reasoning, enhancing logical coherence when incorporating new knowledge. We propose a new KE framework: DeepEdit (Depth-first Search-based Constrained Decoding for Knowledge Editing), which enhances LLMs’s ability to generate coherent reasoning chains with new knowledge through depth-first search. Our search selects the most important knowledge that satisfies our constraints as the reasoning step to efficiently increase the reasoning depth. In addition to DeepEdit, we propose two new KE benchmarks: MQuAKE-2002 and MQuAKE-hard, which provide more precise and challenging assessments of KE approaches. Qualitatively, DeepEdit enables LLMs to produce succinct and coherent reasoning chains involving new knowledge. Quantitatively, it yields significant improvements on multiple KE benchmarks.
@@ 1 Introduction
Knowledge editing (KE) aims to enable large language models (LLMs) to incorporate new or updated knowledge, in contrast to relying solely on outdated parametric knowledge. Multi-hop question answering is the challenging task to evaluate KE Zhong et al. (2023), which requires models to answer questions corresponding to a chain of facts including new knowledge (see Figure 1(a)).
@@
Recent literature has mainly focused on performing KE through in-context editing Zhong et al. (2023); Cohen et al. (2023); Zheng et al. (2023), which puts new knowledge in the input context to fulfill KE (see Figure 1(b)). What new knowledge to use and how to utilize new knowledge is decided by the LLM generators’ implicit reasoning. However, the LLM generator’s hallucinations against the in-context new knowledge can harm the logical coherence of multi-step reasoning and result in incorrect answers Zhang et al. (2023).
To get the correct answers to multi-hop questions with new knowledge, we should ensure that the adjacent reasoning steps are logically coherent, and new knowledge is properly integrated into reasoning (see Fig. Figure 1(a). In this work, we explore using decoding constraints and search algorithms to explicitly control LLMs’ reasoning with new knowledge.
@@
The central idea of this paper is to view KE as a problem of constrained decoding. We design the decoding constraints that facilitate LLMs to soundly incorporate new knowledge in LLMs reasoning. Specifically, we propose the following constraints: Conciseness requires every reasoning step to be identical; Coherence requires the adjacent reasoning steps to be coherent; Receptiveness requires new knowledge to replace the conflicted parametric knowledge; Pertinence requires every reasoning step to be relevant to the target question (see Figure 2). A multi-step reasoning chain that leads to the correct answer should satisfy the above constraints.
@@
We incorporate the above constraints into a new KE framework: DeepEdit (Depth-first Search-based Constrained Decoding for Knowledge Editing). When producing every reasoning step, DeepEdit searches through the parametric and new knowledge to find the knowledge items that satisfy our constraints, termed as valid step candidates. Among the valid candidates, DeepEdit prioritizes more important ones as the new reasoning step for further reasoning. We comprehensively evaluate the importance of different candidates based on their semantic relationships and the prior that the new knowledge is more informative to LLMs than the parametric knowledge.
Note that the classical depth-first search has a backtracking mechanism that enables us to traverse all the reasoning chains of valid reasoning steps Tarjan (1972). However, such backtracking is undoubtedly time-consuming and unnecessary after an answer is found. Therefore, we design an early-stopping mechanism to stop the depth-first search when an answer is found by LLMs. Such an early stopping significantly improves the reasoning efficiency by reducing redundant iterations, as validated by the qualitative analysis (see Section 3.3) and empirical results (see Section 5.4).
The advantages of DeepEdit are two-fold. First, DeepEdit is flexibly applicable to any black-box LLM without requiring access to model parameters or token-wise distributions. Second, DeepEdit improves both the receptiveness of new knowledge and the coherence of multi-step reasoning by directly controlling the LLMs’ reasoning with decoding constraints. Unlike previous KE methods that edit either the model parameters Zhu et al. (2020) or the input prompts Cohen et al. (2023), we directly control the decoding of LLMs to incorporate the new knowledge into outputs, which opens up a new direction for KE in LLMs.
@@
To provide precise and challenging assessments of KE methods, we build two new benchmarks, MQuAKE-2002, and MQuAKE-hard, which resolve the knowledge-conflicting annotation mistakes in the popular KE benchmark MQuAKE-3k Zhong et al. (2023). We evaluate DeepEdit on MQuAKE-3k, MQuAKE-2002, and MQuAKE-hard Zhong et al. (2023); Meng et al. (2022a). Qualitatively, DeepEdit enhances LLMs to produce succinct reasoning with new knowledge, as shown in Figure 9. Quantitatively, DeepEdit improves the KE for popular LLMs by a significant margin OpenAI (2022); Touvron et al. (2023).
@@ 2 Related Work
Previous studies have explored multiple methods for editing the knowledge of LLMs by introducing new knowledge into static model artifacts Zhu et al. (2020); Sotoudeh and Thakur (2019); Dai et al. (2021); Hase et al. (2021); Zhou et al. (2023); Dong et al. (2022); Huang et al. (2023). Some of these approaches involve identifying and modifying model weights associated with specific concepts Meng et al. (2022a, b); Dai et al. (2021), as well as rapid adaptation facilitated by a compact auxiliary editing network Mitchell et al. (2022); De Cao et al. (2021); Mitchell et al. (2021). Given the fast-growing parameter sizes of large language models (LLMs) Zhao et al. (2023), frequently updating LLMs with new knowledge through retraining is more and more expensive. Hence, it is vital to effectively edit the LLMs’ knowledge without retraining. Recent work observes the superior KE performance of in-context editing. MeLLo Zhong et al. (2023) designs a single prompt to alternately conduct the text generation and new knowledge analysis using the LLM generator. Cohen et al. (2023) propose to append the new knowledge the head of the input prompt. In-context editing lacks direct control of the LLMs’ outputs and how the new knowledge influences the outputs is not traceable.
Different from the prior work, we propose a new KE method DeepEdit, which effectively augments LLMs on KE with constrained decoding. We illustrate a comparison between DeepEdit and existing KE methods in Table 1. Our DeepEdit outperforms the prior KE methods by editing the LLMs’ outputs directly, which paves the way to a new direction of LLMs’ KE.
Constrained Decoding
@@ 3 Methodology
Multi-hop question answering with new knowledge assumes a database that stores all new facts, which LLMs should adhere to, and posits that only a few facts can influence the ground-truth answer to each question. In this section, we will introduce how we integrate the decoding constraints into the design of a novel KE method, DeepEdit (Depth-first Search based Decoding for Knowledge Editing), to enhance the KE of black-box LLMs.
@@ 3.1 Constraint Designing for Knowledge Editing
On an LLM’s multi-step reasoning chain to answer a multi-hop question involving new knowledge, we find the following constraints on reasoning steps can help LLMs to give the correct answer:
@@
Among the above four constraints, Conciseness removes loops in the reasoning chain; Coherence guarantees the coherence of reasoning; Receptiveness ensures the LLMs’ receptiveness to new knowledge; Pertinence guides LLMs to answer the target question. We visualize different constraints in Figure 2. The above constraints are easy to follow for LLMs when there is no new knowledge. However, the utilization of in-context new knowledge is influenced by the LLMs’ hallucinations, which can easily break the aforementioned constraints. In this work, we apply the above constraints to incorporate the new knowledge into LLMs’ reasoning.
@@ 3.2 Constraint Verification on Every Step
We term the reasoning steps that satisfy all the constraints as valid steps. Because an invalid reasoning step would mislead LLMs into generating incorrect subsequent steps, we verify the constraints at the decoding of each reasoning step.
At every iteration, we take both the parametric and new knowledge as the step candidates. First, we let LLMs generate one reasoning step with temperature as 00, which represents the parametric knowledge that the LLM believes to be most useful to answer the target question. Second, we retrieve $N$ new knowledge closest to the generated reasoning step. The distance between knowledge and the generated step is measured by the semantic distance of sentence-level embeddings given by a pre-trained BERT model Devlin et al. (2019). The closer the context of the new knowledge is to the generated step, the more likely it is that the new knowledge will be helpful in answering the target question. We take these $N+1$ knowledge as the reasoning step candidates, with examples shown in Figure 3.
We verify the constraints on the aforementioned step candidates to identify the valid ones. In principle, every constraint’s verification can be seen as a binary classification problem on a pair of sentences. For example, Coherence’s verification can be completed by assessing whether a pair of reasoning steps is relevant to each other or not. Therefore, one option to do constraint verification is to train a binary classifier for every constraint. However, such a solution is not data or time-efficient. In this work, we develop four specialized in-context learning-based LLM agents to act as the verifiers. Every verifier’s prompt includes $D$ positive and negative demonstrations randomly sampled from MQuAKE Zhong et al. (2023). In our experiments, we set $D=2$ and $N=2$ by default.
@@ 3.3 Efficient Reasoning with Depth-first Search
At every iteration, there may exist more than valid step candidates. We can consider each valid step candidate simultaneously as the new reasoning step and make further reasoning based on each of them separately. This design can be seen as a breadth-first search of a valid multi-hop reasoning chain. It is undoubtedly time-consuming because the time complexity grows exponentially with the number of reasoning steps. This raises a natural question:
Can we increase the depth of reasoning chains more efficiently than the breadth-first search?
Our answer is yes. We propose prioritizing more important step candidates as the reasoning step for further reasoning. Actually, different valid step candidates are not equally important. First, the edited facts should be more important than the generated step, since the former generally contains new knowledge unknown to LLMs. Secondly, the edited facts that are closer to the generated step are more important. They are more helpful in answering the target question due to their higher relevance to the context generated by the LLM’s reasoning. Based on the above prior, in every iteration, we select the most important knowledge as the next reasoning step and do further reasoning based on it, as shown in Figure 4.
@@
Figure 5: The iteration-wise visualization of our DeepEdit on an instance in MQuAKE-3k. In every iteration, the retrieved new knowledge (highlighted in green) is ranked from top to bottom as step candidates based on their semantic distances to the generated step, in ascending order.
We name our method as DeepEdit: depth-first search-based constrained decoding for knowledge editing. If we follow the classical depth-first search algorithm Tarjan (1972), we should include a backtracking mechanism to traverse all the reasoning chains of valid steps. However, such backtracking is undoubtedly time-consuming and unnecessary after an answer is found. Therefore, we design an early-stopping mechanism to stop the search when an answer is found by LLMs. We visualize DeepEdit in Figure 1 the example iteration-wise outputs in Figure 5.
Assume the number of reasoning hops as $h$, and the number of valid step candidates per iteration as $c$. The time complexity of breadth-first search is $\mathcal{O}(c^{h})$. In contrast, thanks to our depth-first search design with the early-stopping mechanism, our model has a significantly lower time complexity of $\mathcal{O}(h)$ in the best case. Empirical results validate DeepEdit’s superior efficiency compared with the breadth-first search (see Section 5.4).
The strategy of depth-first search has been utilized by some research on LLMs, such as tree-of-thoughts Yao et al. (2023). Different from them, our DeepEdit’s depth-first search is designed to efficiently find the valid multi-hop reasoning chain incorporating new knowledge. Our DeepEdit is free to be applied to all black-box LLMs: it bypasses the requirement of accessing the token-wise distributions because our decoding strategy is defined at the reasoning step level.
@@ 5.1 Experimental Settings
LLM as Constraint Verfier CON COH REC PER
LLaMA2-7B-Chat 82.3 77.9 86.3 78.2
GPT-3.5-Turbo-Instruct 92.8 91.5 92.2 86.7
@@ Appendix B Analysis on the Data where the Edited Facts do not Change the Ground-Truth Answers
Appendix C Case Study
@@
Prompt Text Generation Prompt of DeepEdit: Question: What is the capital of the country where Plainfield Town Hall is located? Thoughts with New Knowledge: Plainfield Town Hall is located in the country of the United States of America.# The capital of United States is Washington, D.C.# Washington, D.C. is the answer. Answer: Washington, D.C. Question: In which country is the company that created Nissan 200SX located? Thoughts with New Knowledge: Nissan 200SX was created by Nissan.# Nissan is located in the country of Japan.# Japan is the answer. Answer: Japan Question: Which continent is the country where the director of "My House Husband: Ikaw Na!" was educated located in? Thoughts with New Knowledge: The director of "My House Husband: Ikaw Na!" is Jose Javier Reyes.# Jose Javier Reyes was educated at De La Salle University.# De La Salle University is located in the country of Philippines.# Philippines is located in the continent if Asia.# Asia is the answer. Answer: Asia
Figure 10: Text Generation Prompt of DeepEdit Zhong et al. (2023)
Prompt Text Generation Prompt of MeLLo: Question: What is the capital city of the country of citizenship of Ivanka Trump’s spouse? Subquestion: Who is Ivanka Trump’s spouse? Generated answer: Ivanka Trump’s spouse is Jared Kushner. Retrieved fact: Ivanka Biden is married to Samantha Cameron. Retrieved fact does not contradict to generated answer, so the intermediate answer is: Jared Kushner Subquestion: What is the country of citizenship of Jared Kushner? Generated answer: The country of citizenship of Jared Kushner is United States. Retrieved fact: Jared Kushner is a citizen of Canada. Retrieved fact contradicts to generated answer, so the intermediate answer is: Canada Subquestion: What is the capital city of Canada? Generated answer: The capital city of Canada is Ottawa. Retrieved fact: The capital city of United States of America is Seattle. Retrieved fact does not contradict to generated answer, so the intermediate answer is: Ottawa Final answer: Ottawa Question: Who is the head of state of the country where Rainn Wilson holds a citizenship? Subquestion: What is the country of citizenship of Rainn Wilson? Generated answer: Rainn Wilson is a citizen of United States of America. Retrieved fact: Rainn Wilson is a citizen of Croatia. Retrieved fact contradicts to generated answer, so the intermediate answer is: Croatia Subquestion: What is the name of the current head of state in Croatia? Generated answer: The head of state of Croatia is President Zoran Milanović. Retrieved fact: The name of the current head of state in Croatia is Kolinda Grabar-Kitarović. Retrieved fact contradicts to generated answer, so the intermediate answer is: Kolinda Grabar-Kitarović Final answer: Kolinda Grabar-Kitarović Question: Who is the spouse of the head of state in United States of America? Subquestion: Who is the head of state in United States of America? Generated answer: The US president is Donald Trump. Retrieved fact: The head of state in United States of America is Joe Biden. Retrieved fact contradicts to generated answer, so the intermediate answer is: Joe Biden Subquestion: Who is the spouse of Joe Biden? Generated answer: The spouse of Joe Biden is Jill Biden. Retrieved fact: The spouse of Joe Bill is Evan Austin. Retrieved fact does not contradict to generated answer, so the intermediate answer is: Jill Biden Final answer: Jill Biden Question: On which continent is the country of citizenship of the founder of the manufacturer of iPhone 5 situated? Subquestion: Which company is iPhone 5 produced by? Generated answer: iPhone 5 is produced by Apple. Retrieved fact: The company that produced iPhone 5 is Iveco. Retrieved fact contradicts to generated answer, so the intermediate answer is: Iveco Subquestion: Who is the founder of Iveco? Generated answer: Iveco was founded by Giovanni Agnelli. Retrieved fact: House of Bonaparte was founded by Gustav I of Sweden. Retrieved fact does not contradict to generated answer, so the intermediate answer is: Giovanni Agnelli Subquestion: What is the country of citizenship of Giovanni Agnelli? Generated answer: Giovanni Agnelli is a citizen of Italy. Retrieved fact: Giovanni Agnelli is a citizen of Niger. Retrieved fact contradicts to generated answer, so the intermediate answer is: Niger. Subquestion: On which continent is Niger situated? Generated answer: Niger is situated on Africa. Retrieved fact: Kingdom of England is located in the continent of North America. Retrieved fact does not contradict to generated answer, so the intermediate answer is: Africa. Final answer: Africa
Figure 11: Text Generation Prompt of MeLLo Zhong et al. (2023).